OpenStudy (anonymous):

how would I put 10^(4x-1)=1000 in logarithmic form?

OpenStudy (campbell_st):

why... all you need to do is write 1000 as a power of 10 \[10^{(4x -1)} = 10^3\] just equate the powers

OpenStudy (unklerhaukus):

\[\log_by=x\quad\iff\quad y=b^x\]

OpenStudy (unklerhaukus):

the base \(b\) will be \(10\)

OpenStudy (anonymous):

So ill have (4x-1)log10^3?

OpenStudy (unklerhaukus):

what happend to your equal sign?

OpenStudy (anonymous):

Sp (4x-1)log10=10^3?

OpenStudy (anonymous):

*So

OpenStudy (campbell_st):

nope you will have \[(4x -1)\log_{10} (10) = 3 \log_{10}(10)\]

OpenStudy (campbell_st):

if you take the log in an equation... its done to both sides.

OpenStudy (campbell_st):

so solving for x its 4x -1 = 3

OpenStudy (anonymous):

Thank you guys

OpenStudy (unklerhaukus):

has a neat solution